OpenStudy (anonymous):

Bette is fencing in an area for her dogs. one side of the rectangularr area is 20 ft long. she wants to fence in the area of 500 sq. ft. How many feet of fencing will bette need? @smileyxl3

OpenStudy (smileyxl3):

Area Formula \[A=L \times W\] \[500=20 \times ?\]

OpenStudy (smileyxl3):

woops hold on

OpenStudy (smileyxl3):

Forgot to do perimeter >.< P = 2(20) + 2(25) P = 40 + 50 P = 90 <- that's your answer...

OpenStudy (smileyxl3):

@cupcakestoner I updated the answer